OpenStudy (anonymous):

please help h4p2 mit 6.002 how find vo in p2

OpenStudy (nurali):

\[vo=\frac{ Vi(R _{L}) }{ R _{L}+R _{i n} }\]

OpenStudy (anonymous):

thanks! but in part second when zener diode place in circuit how to find VO & vo?

OpenStudy (nurali):

\[c-V _{o}=\frac{ V _{1}(R _{L}) }{ R _{L}+R _{i n}}\] \[d-v _{o}=\frac{ V _{i}(R _{L}) }{ R _{L}+R _{i n}}\]

OpenStudy (anonymous):

thanks and last Q RL ?

OpenStudy (anonymous):

plz give me formula for this RL

OpenStudy (nurali):

\[V _{o}=V _{i}*\frac{ R _{L _{\min}} }{ R _{L _{\min}}+R _{IN} }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plz explain RLmin

OpenStudy (anonymous):

sajid have u done h3p4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can u tell me answers of h3p2 and h3p4 d1 and d2 only

OpenStudy (anonymous):

0.000182 D1 0.00024 D2

OpenStudy (anonymous):

It is for h3p4

OpenStudy (anonymous):

but it says wrong

OpenStudy (anonymous):

what is your v1 , v2 and R value?

OpenStudy (anonymous):

v1=2,,v2=3,,r=4.7

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

waiting ....thanks

OpenStudy (anonymous):

check this D1= 0.74 D2= 1.59

OpenStudy (anonymous):

sorry bro still wrong

OpenStudy (anonymous):

WHAT IS your peak voltage

OpenStudy (anonymous):

10

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i put it 7.5 b/c it is mine

OpenStudy (anonymous):

okayy...

OpenStudy (anonymous):

now check this D1=1.277

OpenStudy (anonymous):

sorry still wrong

OpenStudy (anonymous):

d2=0.85

OpenStudy (anonymous):

1 min

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

D1=0.0012 it will be correct i am sure

OpenStudy (anonymous):

check it

OpenStudy (anonymous):

really sorry bro still wrong

OpenStudy (anonymous):

r u sure abt ur v1,v2,r and vs values?

OpenStudy (anonymous):

yes v1=2...v2=3...r=4.7... What is the maximum positive voltage (in Volts) that can appear at v? ans=2

OpenStudy (anonymous):

What is the minimum (negative) voltage (in Volts) that can appear at v? ans=-3 correct

OpenStudy (anonymous):

and peak vlatage i.e vs ??

OpenStudy (anonymous):

10.0

OpenStudy (anonymous):

and R = 4.7K ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

D1=0.00127 BRO this the answer if still u have problem then i will give u formula solve it for urself

OpenStudy (anonymous):

great right answer and d2 <give me formula> thanks alot really thanks

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

d2= 0.000851

OpenStudy (anonymous):

bingo great thanks alot thanks...

OpenStudy (anonymous):

now plz clik on best response

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

bye thanks again

OpenStudy (anonymous):

u welcom